OpenStudy (anonymous):

Identify all of the following solutions of square root of x plus 10 end root plus 2 equals x.

OpenStudy (anonymous):

sqrt(x + 10) = x - 2 Squaring both sides: x + 10 = (x - 2)2 = x2 - 4x +

OpenStudy (anonymous):

Are you there